OpenStudy (anonymous):

\[ \lim_{x \rightarrow \infty} 3/x\]

OpenStudy (lalaly):

\[= 3\lim_{x \rightarrow \infty} \frac{1}{x} = 3 \times 0 = 0\]

OpenStudy (anonymous):

when the limit is indefinite how 1/3's value can be equal to zero?? i kno that its equal but never searched why it is :D